# ReCast: Reliability-aware Codebook Assisted Lightweight Time Series Forecasting

### v1

@@ 1 Introduction
In recent years, time series forecasting has gained significant attention due to its critical applications in various real-world applications, including finance, energy, healthcare, and industrial automation (wen2023transformers; ma2023dynamic; qiu2024tfb; shibo2025hdt). Capturing complex and irregular temporal patterns accurately remains a primary challenge in this domain. Conventional approaches typically address this complexity by globally decomposing time series into trend, seasonal, and residual components, and modeling these components independently (Autoformer; Fedformer; hu2025adaptive). However, while effective for structured or periodic data, such global decomposition methods often underperform when faced with dynamic and noisy real-world time series (tang2025unlocking). Moreover, these methods typically involve considerable model complexity, which limits their practicality in resource-limited environments (ansarichronos).
To address these challenges, we introduce a novel REliability-aware Codebook-ASsisted Time series forecasting (ReCast) framework, focusing explicitly on capturing local patterns. Observing that many real-world series exhibit recurring local shapes rather than clear global regularities (yeh2016matrix), ReCast quantizes these local shapes into a learnable codebook, generating discrete embeddings to represent evolving patterns. This codebook-based representation not only captures salient local structures but also reduces model complexity, enabling a inherently lightweight forecasting design. Meanwhile, residual modeling is introduced to capture irregular variations not adequately represented by the quantized embeddings, ensuring robustness to fluctuations without excessively increasing model size.
@@
More importantly, it can be observed that the performance of ReCast strongly depends on the stability and adaptability of the codebook (guo2023compact). Therefore, we propose an incremental codebook update mechanism centered on a reliability-aware scoring method. This method can robustly guide the update process in response to evolving data distributions, striking a balance between stability and adaptability. Our contributions include:
@@ 2.1 Deep Learning-based Forecasting
Recent advances in deep learning have significantly improved time series forecasting by leveraging powerful representation learning capabilities. The convolutional neural network(CNN)-based approaches (wu2023timesnet) introduce local receptive fields to capture short-term dynamics efficiently, but they lack the ability to capture long-range dependencies. Transformer-based models (liu2023itransformer; nietime) address this issue by employing self-attention to model global temporal interactions, achieving strong performance across benchmarks. Nonetheless, the quadratic complexity of attention and sensitivity to noise restrict their scalability and robustness in real-world scenarios. In parallel, lightweight MLP-based architectures (tang2025unlocking; ma2024u) have recently emerged as promising alternatives, offering high efficiency but often struggling to represent heterogeneous and irregular patterns effectively.
@@ 2.2 Patch-based Representation Learning
To improve efficiency and capture fine-grained structures, patch-based strategies have gained increasing attention in time series modeling. Instead of processing sequences at raw temporal resolution, some methods (wang2025towards; nietime; tang2025unlocking) divide time series into non-overlapping or partially overlapping patches, enabling models to operate on compact representations and reduce sequence length. While effective in long-horizon forecasting, these methods typically rely on continuous embeddings without explicit mechanisms to leverage recurring local shapes, which are prevalent in real-world time series. Vector quantization (VQ) (van2017neural) provides a complementary perspective by discretizing local segments into a finite set of codewords, facilitating representation reuse and improving robustness, as extensively explored in domains such as vision and speech (tian2024visual; wu2025janus). Recent attempts (shibo2025hdt; ansarichronos) to integrate quantization into time series tasks demonstrate its potential to capture recurring patterns efficiently. However, static or heuristic codebooks fail to adapt to real-world data dynamics.
@@ 3.1 Patch-wise Quantization
where $q_{i}\in\{1,\cdots,K\}$ is the discrete index associated with patch $\textbf{p}_{i}$. $K$ is the number of codewords. To reduce computational cost and suppress redundant local fluctuations, we apply downsampling $Dsamp(\cdot)$ on patches prior to quantization. This is supported by the well-established assumption in time series modeling that local patterns demonstrate invariance across scales and redundant morphology (lu2022matrix), which makes resolution reduction both meaningful and robust (senin2013sax). $\tilde{\textbf{p}}_{i}$ denotes the downsampled patch of $\textbf{p}_{i}$. This operation reduces the dimension of patches to $L_{p}/2$, significant savings in codebook matching, storage, and embedding projection. Additionally, it helps the codebook focus on salient structures, improving robustness and generalization to noisy or distribution shifts. The discrete embeddings for the full input series is organized as $\textbf{Q}_{x}=[\textbf{Q}_{1};\cdots;\textbf{Q}_{C}]$, and $\textbf{Q}_{i}=\{q_{j}\}_{j=(i-1)\cdot N+1}^{i\cdot N}$ represents the discrete embedding of $i$-th variable. This discrete embeddings serves as the input to downstream forecasting modules.
@@ 3.3 Codebook Construction and Updating
At each epoch, we cluster the patches and obtain cluster centers. These centers are the representative local patterns that can be used to construct pseudo codebooks in the current epoch. The clustering objects are randomly sampled patches from the input. This random sampling reduces computational cost and prevents overfitting (lu2022matrix; senin2013sax). To ensure the efficiency, we express the energy function $\mathcal{L}_{c}$ of clustering in the form of matrix operation:
@@ 3.4 Reliability-aware Scoring
The $\textbf{w}_{je}^{t}$ measures the OOD sensitivity of $\hat{\textbf{S}}^{t}$ by capturing potentially novel or rare patterns, estimated from assignment frequency and variance. The function is similar to joint-energy (duvenaud2020your):
@@
where $w_{je,k}^{t}\in\textbf{w}_{je}^{t}$ is the weight assigned to the $k$-th cluster center. Higher value of $w_{je,k}^{t}$ indicates lower selection probabilities for the $k$-th cluster. By increasing its corresponding weight, we can prevent the embedding space of the codebook from collapsing into a few fixed codewords, and evaluate adaptability to OOD data (duvenaud2020your).
@@
In ReCast, each pseudo codeword is associated with three normalized reliability scores: $\textbf{w}_{rep}^{t}$, $\textbf{w}_{\Delta}^{t}$, and $\textbf{w}_{je}^{t}$. While these metrics are complementary, their relative importance may vary across epochs and data regimes. Directly assigning fixed weights can be suboptimal or unstable, especially when some metrics are noisy or biased due to transient data conditions (duchi2019variance). Thus, we formulate the fusion of reliability metrics as a distributionally robust optimization (DRO) problem (qi2021online). The goal is to obtain a conservative estimate of a codeword’s reliability by minimizing the expected reliability under the worst-case weighting distribution over the three metrics.
@@ 4.1 Datasets and Baselines
We evaluate the proposed ReCast on 8 widely used real-world datasets: Electricity (ECL), Traffic, Weather, Solar (liu2022scinet; Autoformer), and 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2) (zhou2021informer). The full details of datasets presented in Appendix B.1. To evaluate the performance of ReCast, we compare it against 8 representative SOTA models from recent years: Transformer-based models: TQNet (lintemporal), iTransformer (liu2023itransformer), PatchTST (nietime); CNN-based model: TimeNet (wu2023timesnet); MLP-based models: PatchMLP (kong2025unlocking), CycleNet (lin2024cyclenet), DLinear (Zeng2022AreTE).
@@ 4.2 Metrics and Implementation Details
The models are evaluated based on both Mean Squared Error (MSE) and Mean Absolute Error (MAE). ReCast is implemented using Pytorch (paszke2019pytorch) and trained on an Nvidia L40 GPU (48GB). The detailed implementations are described in Appendix B.2. The corresponding pseudocode of ReCast is provided in Appendix B.3.

### v2

@@ 1 Introduction
In recent years, time series forecasting has gained significant attention due to its critical applications in various real-world applications, including finance, energy, healthcare, and industrial automation (Wen et al. 2023; Ma et al. 2023; Qiu et al. 2024; Shibo et al. 2025). Capturing complex and irregular temporal patterns accurately remains a primary challenge in this domain. Conventional approaches typically address this complexity by globally decomposing time series into trend, seasonal, and residual components, and modeling these components independently (Wu et al. 2021; Zhou et al. 2022; Hu et al. 2025). However, while effective for structured or periodic data, such global decomposition methods often underperform when faced with dynamic and noisy real-world time series (Tang and Zhang 2025). Moreover, these methods typically involve considerable model complexity, which limits their practicality in resource-limited environments (Ansari et al. 2025).
To address these challenges, we introduce a novel REliability-aware Codebook-ASsisted Time series forecasting (ReCast) framework, focusing explicitly on capturing local patterns. Observing that many real-world series exhibit recurring local shapes rather than clear global regularities (Yeh et al. 2016), ReCast quantizes these local shapes into a learnable codebook, generating discrete embeddings to represent evolving patterns. This codebook-based representation not only captures salient local structures but also reduces model complexity, enabling a inherently lightweight forecasting design. Meanwhile, residual modeling is introduced to capture irregular variations not adequately represented by the quantized embeddings, ensuring robustness to fluctuations without excessively increasing model size.
@@
More importantly, it can be observed that the performance of ReCast strongly depends on the stability and adaptability of the codebook (Guo et al. 2023). Therefore, we propose an incremental codebook update mechanism centered on a reliability-aware scoring method. This method can robustly guide the update process in response to evolving data distributions, striking a balance between stability and adaptability. Our contributions include:
@@ 2.1 Deep Learning-based Forecasting
Recent advances in deep learning have significantly improved time series forecasting by leveraging powerful representation learning capabilities. The convolutional neural network(CNN)-based approaches (Wu et al. 2023) introduce local receptive fields to capture short-term dynamics efficiently, but they lack the ability to capture long-range dependencies. Transformer-based models (Liu et al. 2023; Nie et al. 2023) address this issue by employing self-attention to model global temporal interactions, achieving strong performance across benchmarks. Nonetheless, the quadratic complexity of attention and sensitivity to noise restrict their scalability and robustness in real-world scenarios. In parallel, lightweight MLP-based architectures (Tang and Zhang 2025; Ma et al. 2024) have recently emerged as promising alternatives, offering high efficiency but often struggling to represent heterogeneous and irregular patterns effectively.
@@ 2.2 Patch-based Representation Learning
To improve efficiency and capture fine-grained structures, patch-based strategies have gained increasing attention in time series modeling. Instead of processing sequences at raw temporal resolution, some methods (Wang et al. 2025; Nie et al. 2023; Tang and Zhang 2025) divide time series into non-overlapping or partially overlapping patches, enabling models to operate on compact representations and reduce sequence length. While effective in long-horizon forecasting, these methods typically rely on continuous embeddings without explicit mechanisms to leverage recurring local shapes, which are prevalent in real-world time series. Vector quantization (VQ) (Van Den Oord et al. 2017) provides a complementary perspective by discretizing local segments into a finite set of codewords, facilitating representation reuse and improving robustness, as extensively explored in domains such as vision and speech (Tian et al. 2024; Wu et al. 2025). Recent attempts (Shibo et al. 2025; Ansari et al. 2025) to integrate quantization into time series tasks demonstrate its potential to capture recurring patterns efficiently. However, static or heuristic codebooks fail to adapt to real-world data dynamics.
@@ 3.1 Patch-wise Quantization
where $q_{i}\in\{1,\cdots,K\}$ is the discrete index associated with patch $\textbf{p}_{i}$. $K$ is the number of codewords. To reduce computational cost and suppress redundant local fluctuations, we apply downsampling $Dsamp(\cdot)$ on patches prior to quantization. This is supported by the well-established assumption in time series modeling that local patterns demonstrate invariance across scales and redundant morphology (Lu et al. 2022), which makes resolution reduction both meaningful and robust (Senin and Malinchik 2013). $\tilde{\textbf{p}}_{i}$ denotes the downsampled patch of $\textbf{p}_{i}$. This operation reduces the dimension of patches to $L_{p}/2$, significant savings in codebook matching, storage, and embedding projection. Additionally, it helps the codebook focus on salient structures, improving robustness and generalization to noisy or distribution shifts. The discrete embeddings for the full input series is organized as $\textbf{Q}_{x}=[\textbf{Q}_{1};\cdots;\textbf{Q}_{C}]$, and $\textbf{Q}_{i}=\{q_{j}\}_{j=(i-1)\cdot N+1}^{i\cdot N}$ represents the discrete embedding of $i$-th variable. This discrete embeddings serves as the input to downstream forecasting modules.
@@ 3.3 Codebook Construction and Updating
At each epoch, we cluster the patches and obtain cluster centers. These centers are the representative local patterns that can be used to construct pseudo codebooks in the current epoch. The clustering objects are randomly sampled patches from the input. This random sampling reduces computational cost and prevents overfitting (Lu et al. 2022; Senin and Malinchik 2013). To ensure the efficiency, we express the energy function $\mathcal{L}_{c}$ of clustering in the form of matrix operation:
@@ 3.4 Reliability-aware Scoring
The $\textbf{w}_{je}^{t}$ measures the OOD sensitivity of $\hat{\textbf{S}}^{t}$ by capturing potentially novel or rare patterns, estimated from assignment frequency and variance. The function is similar to joint-energy (Duvenaud et al. 2020):
@@
where $w_{je,k}^{t}\in\textbf{w}_{je}^{t}$ is the weight assigned to the $k$-th cluster center. Higher value of $w_{je,k}^{t}$ indicates lower selection probabilities for the $k$-th cluster. By increasing its corresponding weight, we can prevent the embedding space of the codebook from collapsing into a few fixed codewords, and evaluate adaptability to OOD data (Duvenaud et al. 2020).
@@
In ReCast, each pseudo codeword is associated with three normalized reliability scores: $\textbf{w}_{rep}^{t}$, $\textbf{w}_{\Delta}^{t}$, and $\textbf{w}_{je}^{t}$. While these metrics are complementary, their relative importance may vary across epochs and data regimes. Directly assigning fixed weights can be suboptimal or unstable, especially when some metrics are noisy or biased due to transient data conditions (Duchi and Namkoong 2019). Thus, we formulate the fusion of reliability metrics as a distributionally robust optimization (DRO) problem (Qi et al. 2021). The goal is to obtain a conservative estimate of a codeword’s reliability by minimizing the expected reliability under the worst-case weighting distribution over the three metrics.
@@ 4.1 Datasets and Baselines
We evaluate the proposed ReCast on 8 widely used real-world datasets: Electricity (ECL), Traffic, Weather, Solar (Liu et al. 2022; Wu et al. 2021), and 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2) (Zhou et al. 2021). The full details of datasets presented in Appendix B.1. To evaluate the performance of ReCast, we compare it against 8 representative SOTA models from recent years: Transformer-based models: TQNet (Lin et al. 2025), iTransformer (Liu et al. 2023), PatchTST (Nie et al. 2023); CNN-based model: TimeNet (Wu et al. 2023); MLP-based models: PatchMLP (Kong et al. 2025), CycleNet (Lin et al. 2024), DLinear (Zeng et al. 2023).
@@ 4.2 Metrics and Implementation Details
The models are evaluated based on both Mean Squared Error (MSE) and Mean Absolute Error (MAE). ReCast is implemented using Pytorch (Paszke et al. 2019) and trained on an Nvidia L40 GPU (48GB). The detailed implementations are described in Appendix B.2. The corresponding pseudocode of ReCast is provided in Appendix B.3.